# Yifan Wang1∗ Jianjun Zhou123∗ Haoyi Zhu1 Wenzheng Chang1 Yang Zhou1

### v1

@@ Abstract
We introduce $\pi^{3}$, a feed-forward neural network that offers a novel approach to visual geometry reconstruction, breaking the reliance on a conventional fixed reference view. Previous methods often anchor their reconstructions to a designated viewpoint, an inductive bias that can lead to instability and failures if the reference is suboptimal. In contrast, $\pi^{3}$ employs a fully permutation-equivariant architecture to predict affine-invariant camera poses and scale-invariant local point maps without any reference frames. This design makes our model inherently robust to input ordering and highly scalable. These advantages enable our simple and bias-free approach to achieve state-of-the-art performance on a wide range of tasks, including camera pose estimation, monocular/video depth estimation, and dense point map reconstruction. Code and models are publicly available.
@@ 1 Introduction
“Learning always beats inductive biases.” — Jonathan T. Barron, Radiance Fields and the Future of Generative Media
@@
Figure 2: $\pi^{3}$ demonstrates significant convergence acceleration (a) and enhanced scalability (b) compared to its non-equivariant counterpart. PI denotes the Permutation-EquIvariant property.
Despite these advances, a critical limitation persists in both classical and modern approaches: the reliance on selecting a single, fixed reference view. The camera coordinate system of this chosen view is treated as the global frame of reference, a practice inherited from traditional Structure-from-Motion (SfM) [11, 4, 24, 20] or Multi-view Stereo (MVS) [9, 25]. We contend that this design choice introduces an unnecessary inductive bias that fundamentally constrains the performance and robustness of feed-forward neural networks. As we demonstrate empirically, this reliance on an arbitrary reference makes existing methods, including the state-of-the-art (SOTA) VGGT [34], highly sensitive to the initial view selection. A poor choice can lead to a dramatic degradation in reconstruction quality, hindering the development of robust and scalable systems (Fig. 3).
To overcome this limitation, we introduce $\pi^{3}$, a robust, scalable, and fully permutation-equivariant method that eliminates reference view-based biases in visual geometry learning. $\pi^{3}$ accepts varied inputs—including single images, video sequences, or unordered image sets from static or dynamic scenes—without designating a reference view. Instead, our model predicts an affine-invariant camera pose and a scale-invariant local pointmap for each frame, all relative to that frame’s own camera coordinate system. By eschewing order-dependent components like frame index positional embeddings and employing a transformer architecture that alternates between view-wise and global self-attention (similar to [34]), $\pi^{3}$ achieves true permutation equivariance. This guarantees a consistent one-to-one mapping between visual inputs and the reconstructed geometry, making the model inherently robust to input order and immune to the reference view selection problem.
@@
Our design yields significant advantages. First, it demonstrates remarkable scalability, with performance consistently improving as the model size increases (Fig. 2 and Fig. 9). Second, it achieves significantly faster training convergence (Fig. 2 and Fig. 9). Finally, our model is substantially more robust; in contrast to previous methods, it shows minimal performance degradation and low standard deviation when the reference frame is changed (Fig. 3 and Table 7).
@@
  • We show that our approach is not only substantially more robust to input view ordering and more scalable with model size, but also converges significantly faster during training.
@@ 2.2 Feed-Forward 3D Reconstruction
Subsequent work has focused on overcoming this limitation. Fast3R [42] represents a significant advance by enabling simultaneous inference on thousands of images, thereby eliminating the need for a costly and fragile global alignment stage. Other approaches have explored simplifying the learning problem itself. For instance, FLARE [47] decomposes the task by first predicting camera poses and then estimating the scene geometry. The current state-of-the-art, VGGT [34], leverages multi-task learning and large-scale datasets to achieve superior accuracy and performance.
@@ 3.1 Permutation-Equivariant Architecture
This property guarantees a consistent one-to-one correspondence between each image and its respective output (e.g., geometry or pose). This design offers several key advantages. First, reconstruction quality becomes independent of the reference view selection, in contrast to prior methods that suffer from performance degradation when the reference view changes. Second, the model becomes more robust to uncertain or noisy observations. Finally, this architecture exhibits superior scalability. These claims are empirically validated in Section 4.
@@ 3.2 Scale-Invariant Local Geometry
Consequently, the training process requires aligning the predicted point maps, $(\hat{\mathbf{X}}_{1},\dots,\hat{\mathbf{X}}_{N})$, with the corresponding ground-truth (GT) set, $(\mathbf{X}_{1},\dots,\mathbf{X}_{N})$. This alignment is accomplished by solving for a single optimal scale factor, $\hat{s}$, which minimizes the depth-weighted L1 distance across the entire image sequence. The optimization problem is formulated as:
[…]
@@
Finally, the point cloud reconstruction loss, $\mathcal{L}_{\text{points}}$, is defined using the optimal scale factor $\hat{s}$:
$\mathcal{L}_{\text{points}}=\sum_{i=1}^{N}\sum_{j=1}^{H\times W}\frac{1}{z_{i,j}}\|\hat{s}\hat{\mathbf{x}}_{i,j}-\mathbf{x}_{i,j}\|_{1}$ (5)
@@
We supervise the predicted confidence map $\mathbf{C}_{i}$ using a Binary Cross-Entropy (BCE) loss, denoted $\mathcal{L}_{\text{conf}}$. The ground-truth target for each point is set to 1 if its L1 reconstruction error, $\frac{1}{z_{i,j}}\|\hat{s}\hat{\mathbf{x}}_{i,j}-\mathbf{x}_{i,j}\|_{1}$, is below a threshold $\epsilon$, and 0 otherwise.
@@ 3.3 Affine-Invariant Camera Pose
Each predicted relative pose $\hat{\mathbf{T}}_{i\leftarrow j}$ is composed of a rotation $\hat{\mathbf{R}}_{i\leftarrow j}\in SO(3)$ and a translation $\hat{\mathbf{t}}_{i\leftarrow j}\in\mathbb{R}^{3}$. While the relative rotation is invariant to this global transformation, the relative translation’s magnitude is ambiguous. We resolve this by leveraging the optimal scale factor, $\hat{s}$, that is computed by aligning the predicted point map to the ground truth (as detailed in a previous section). This single, consistent scale factor is used to rectify all predicted camera translations, allowing us to directly supervise both the rotation and the correctly-scaled translation components.
@@
$\mathcal{L}_{\text{cam}}=\frac{1}{N(N-1)}\sum_{i\neq j}\mathopen{}\mathclose{{}\left(\mathcal{L}_{\text{rot}}(i,j)+\lambda\mathcal{L}_{\text{trans}}(i,j)}\right)$ (8)
@@ 4 Experiments
We evaluate our method on four tasks: camera pose estimation (Sec. 4.1), point map estimation (Sec. 4.2), video depth estimation (Sec. 4.3) and monocular depth estimation (Sec. 4.4). Across all tasks, our method achieves state-of-the-art(SOTA) or comparable performance against existing feed-forward 3D reconstruction methods. To validate the effectiveness of our design, We also conduct several analyses: (1) a robustness evaluation against input image sequence permutations (Sec. 4.5), (2) an ablation study on scale-invariant point maps and affine-invariant camera poses (Sec. 4.6), and (3) a scalability analysis with respect to model size (Sec. 4.7).
@@ 4.1 Camera Pose Estimation
Fast3R [42] sparse 0.096 0.065 0.145 0.093 0.672 0.760 0.135 0.091 0.163 0.104 0.759 0.877
CUT3R [36] 0.094 0.051 0.101 0.050 0.703 0.804 0.104 0.041 0.079 0.031 0.822 0.968
FLARE [47] 0.085 0.058 0.142 0.104 0.695 0.779 0.053 0.024 0.051 0.025 0.877 0.988
VGGT [34] 0.046 0.026 0.057 0.034 0.728 0.842 0.051 0.029 0.066 0.038 0.890 0.981
$\pi^{3}$ (Ours) 0.048 0.028 0.072 0.047 0.742 0.842 0.026 0.015 0.028 0.014 0.916 0.992
Fast3R [42] dense 0.038 0.015 0.056 0.018 0.645 0.725 0.072 0.030 0.050 0.016 0.790 0.934
CUT3R [36] 0.022 0.010 0.027 0.009 0.668 0.762 0.086 0.037 0.048 0.017 0.800 0.953
FLARE [47] 0.018 0.007 0.027 0.014 0.681 0.781 0.023 0.011 0.018 0.008 0.882 0.986
VGGT [34] 0.022 0.008 0.027 0.013 0.663 0.757 0.017 0.010 0.015 0.005 0.893 0.988
$\pi^{3}$ (Ours) 0.015 0.007 0.022 0.011 0.687 0.790 0.015 0.008 0.013 0.005 0.898 0.987
@@ 4.3 Video Depth Estimation
Abs Rel $\downarrow$ $\delta<1.25$uparrow Abs Rel $\downarrow$ $\delta<1.25$uparrow Abs Rel $\downarrow$ $\delta<1.25$uparrow
@@ 4.6 Ablation Study
To make a fair comparison, we use a modified version of VGGT [34], with the tracking and global point map heads disabled, as our Baseline model (1) [Baseline]. We sequentially integrated the Scale-Invariant Pointmap and the Affine-Invariant Camera Pose, resulting in model (2) and model (3) [Full Model], respectively.
Model (1) predicts camera and depth similar to VGGT. Both models (1) and (2) utilize an additional token to designate a reference frame and predict camera poses within that frame’s coordinate system, while model (3) is fully permutation-equivariant.
All models are trained for 80 epochs on images with a resolution of $224\times 224$. The reported metric is the mean of Accuracy and Completeness, averaged across the 7-Scenes [27] and NRGBD [1] datasets. The camera loss for the baseline model remains the same as in VGGT. The detailed architectures of Model (2) and Model (3) can be found in Appendix A.1. All models in the ablation study are trained from scratch, except for the encoder, which is initialized with DINOv2 [18] weights.
As summarized in Table 8, the addition of the scale-invariant pointmap improves the metric from 0.1498 to 0.1145. Our full model, which also incorporates the affine-invariant pose, achieves a significantly better score of 0.0625. This confirms that both components are essential and contribute substantially to the final accuracy and robustness of our architecture.
@@
Table 8: Ablation study on the key components of our model. We show how the performance metric (lower is better) improves as each component is added to the baseline.
ID Model Configuration Metric
(1) Baseline 0.1498
(2) Baseline + Scale-Invariant Pointmap 0.1145
(3) Model (2) + Affine-Invariant Camera Pose (Full Model) 0.0625
4.7 Scalability and Performance Evaluation
Building on the experimental setup described in Section 4.6, we further evaluate the scalability of our fully permutation-equivariant (PI) architecture. We compare it against the non-PI baseline (model (2)) across three model sizes: Small (196.49M parameters), Base (390.13M), and Large (892.37M). The Small model employs a ViT-S encoder, with alternating attention modules configured with a 384-dimensional hidden state, 6 heads, and 24 layers. The Base model uses a ViT-B encoder alongside attention modules with a 768-dimensional hidden state, 12 heads, and 24 layers. The Large model adopts a ViT-L encoder, with a 1024-dimensional hidden state, 16 heads, and 36 layers. As shown in Figure 2, our PI architecture demonstrates clear and consistent advantages. It not only achieves superior final performance with the Large model, achieving a 45% improvement over its baseline, but also exhibits significantly faster convergence, reflecting enhanced sample efficiency. Importantly, these benefits hold across all three model scales, confirming the strong scalability of our approach. Additional experiments are provided in Appendix A.4.
@@ 5 Conclusion
In this work, we introduced $\pi^{3}$, a feed-forward neural network that presents a new paradigm for visual geometry reconstruction by eliminating the reliance on a fixed reference view. By leveraging a fully permutation-equivariant architecture, our model is inherently robust to input ordering and highly scalable. This design choice removes a critical inductive bias found in previous methods, allowing our simple yet powerful approach to achieve state-of-the-art performance on a wide array of tasks, including camera pose estimation, depth estimation, and dense reconstruction. $\pi^{3}$ demonstrates that reference-free systems are not only viable but can lead to more stable and versatile 3D vision models.
@@ A.2 Training Details
We train $\pi^{3}$ in two stages, a process similar to Dust3R [39]. First, the model is trained on a low resolution of 224 $\times$ 224 pixels. Then, it is fine-tuned on images of random resolutions where the total pixel count is between 100,000 and 255,000 and the aspect ratio is sampled from the range [0.5, 2.0], a strategy similar to MoGe [37]. We use a dynamic batch sizing strategy similar to VGGT. In the first stage, we sample 64 images per GPU, and in the second stage, we sample 48 images per GPU. Each batch is composed of 2 to 24 images. Each training stage runs for 100 epochs, with each epoch comprising 1,000 iterations. Our final model is not trained from scratch. Instead, we initialize the weights for the encoder and the alternating attention module from the pre-trained VGGT model, and we keep the encoder frozen during training. For each stage, we use the Adam optimizer with a learning rate that starts at 1 $\times$ $10^{-4}$ and then decays. We train the first stage on 16 A100 GPUs and the second stage on 64 A100 GPUs.

### v2

@@ Abstract
We introduce $\pi^{3}$, a feed-forward neural network that offers a novel approach to visual geometry reconstruction, breaking the reliance on a conventional fixed reference view. Previous methods often anchor their reconstructions to a designated viewpoint, an inductive bias that can lead to instability and failures if the reference is suboptimal. In contrast, $\pi^{3}$ employs a fully permutation-equivariant architecture to predict affine-invariant camera poses and scale-invariant local point maps without any reference frames. This design not only makes our model inherently robust to input ordering, but also leads to higher accuracy and performance. These advantages enable our simple and bias-free approach to achieve state-of-the-art performance on a wide range of tasks, including camera pose estimation, monocular/video depth estimation, and dense point map reconstruction. Code and models are publicly available.
@@ 1 Introduction
Despite these advances, a critical limitation persists in both classical and modern approaches: the reliance on selecting a single, fixed reference view. The camera coordinate system of this chosen view is treated as the global frame of reference, a practice inherited from traditional Structure-from-Motion (SfM) [11, 4, 24, 20] or Multi-view Stereo (MVS) [9, 25]. We contend that this design choice introduces an unnecessary inductive bias that fundamentally constrains the performance and robustness of feed-forward neural networks. As we demonstrate empirically, this reliance on an arbitrary reference makes existing methods, including the state-of-the-art (SOTA) VGGT [34], highly sensitive to the initial view selection. A poor choice can lead to a dramatic degradation in reconstruction quality, hindering the development of robust systems (Fig. 2).
To overcome this limitation, we introduce $\pi^{3}$, a robust, accurate, and fully permutation-equivariant method that eliminates reference view-based biases in visual geometry learning. $\pi^{3}$ accepts varied inputs—including single images, video sequences, or unordered image sets from static or dynamic scenes—without designating a reference view. Instead, our model predicts an affine-invariant camera pose and a scale-invariant local pointmap, with the pointmap being defined in that frame’s own camera coordinate system. By eschewing order-dependent components like frame index positional embeddings and employing a transformer architecture that alternates between view-wise and global self-attention (similar to [34]), $\pi^{3}$ achieves true permutation equivariance. This guarantees a consistent one-to-one mapping between visual inputs and the reconstructed geometry, making the model inherently robust to input order and immune to the reference view selection problem (Table 8).
@@
Our design yields significant advantages. Primarily, it is substantially more robust. Unlike previous methods, our approach demonstrates minimal performance degradation and a low standard deviation when the reference frame is altered (Fig. 2 and Table 7). Furthermore, it enhances reconstruction accuracy over earlier methods that rely on a reference view.
@@ 2.2 Feed-Forward 3D Reconstruction
Subsequent work has focused on overcoming this limitation. Fast3R [42] represents a significant advance by enabling simultaneous inference on thousands of images, thereby eliminating the need for a costly and fragile global alignment stage. Other approaches have explored simplifying the learning problem itself. For instance, FLARE [47] decomposes the task by first predicting camera poses and then estimating the scene geometry. VGGT [34] leverages multi-task learning and large-scale datasets to achieve superior accuracy and performance.
@@ 3.1 Permutation-Equivariant Architecture
This property guarantees a consistent one-to-one correspondence between each image and its respective output (e.g., geometry or pose). This design offers several key advantages. First, reconstruction quality becomes independent of the reference view selection, in contrast to prior methods that suffer from performance degradation when the reference view changes. Second, the model becomes more robust to uncertain or noisy observations. These claims are empirically validated in Section 4.
@@ 3.2 Scale-Invariant Local Geometry
Consequently, the training process requires aligning the predicted point maps, $(\hat{\mathbf{X}}_{1},\dots,\hat{\mathbf{X}}_{N})$, with the corresponding ground-truth (GT) set, $(\mathbf{X}_{1},\dots,\mathbf{X}_{N})$. This alignment is accomplished by solving for a single optimal scale factor, $s^{*}$, which minimizes the depth-weighted L1 distance across the entire image sequence. The optimization problem is formulated as:
$s^{*}=\underset{s}{\arg\min}\sum_{i=1}^{N}\sum_{j=1}^{H\times W}\frac{1}{z_{i,j}}\|s\hat{\mathbf{x}}_{i,j}-\mathbf{x}_{i,j}\|_{1}$ (4)
@@
Finally, the point cloud reconstruction loss, $\mathcal{L}_{\text{points}}$, is defined using the optimal scale factor $s^{*}$:
$\mathcal{L}_{\text{points}}=\frac{1}{3NHW}\sum_{i=1}^{N}\sum_{j=1}^{H\times W}\frac{1}{z_{i,j}}\|s^{*}\hat{\mathbf{x}}_{i,j}-\mathbf{x}_{i,j}\|_{1}$ (5)
@@
We supervise the predicted confidence map $\mathbf{C}_{i}$ using a Binary Cross-Entropy (BCE) loss, denoted $\mathcal{L}_{\text{conf}}$. The ground-truth target for each point is set to 1 if its L1 reconstruction error, $\frac{1}{z_{i,j}}\|s^{*}\hat{\mathbf{x}}_{i,j}-\mathbf{x}_{i,j}\|_{1}$, is below a threshold $\epsilon$, and 0 otherwise.
@@ 3.3 Affine-Invariant Camera Pose
Each predicted relative pose $\hat{\mathbf{T}}_{i\leftarrow j}$ is composed of a rotation $\hat{\mathbf{R}}_{i\leftarrow j}\in SO(3)$ and a translation $\hat{\mathbf{t}}_{i\leftarrow j}\in\mathbb{R}^{3}$. While the relative rotation is invariant to this global transformation, the relative translation’s magnitude is ambiguous. We resolve this by leveraging the optimal scale factor, $s^{*}$, that is computed by aligning the predicted point map to the ground truth (as detailed in a previous section). This single, consistent scale factor is used to rectify all predicted camera translations, allowing us to directly supervise both the rotation and the correctly-scaled translation components.
@@
$\mathcal{L}_{\text{cam}}=\frac{1}{N(N-1)}\sum_{i\neq j}\mathopen{}\mathclose{{\left(\mathcal{L}_{\text{rot}}(i,j)+\lambda_{trans}\mathcal{L}_{\text{trans}}(i,j)}}\right)$ (8)
@@ 4 Experiments
We evaluate our method on four tasks: camera pose estimation (Sec. 4.1), point map estimation (Sec. 4.2), video depth estimation (Sec. 4.3) and monocular depth estimation (Sec. 4.4). Across all tasks, our method achieves state-of-the-art(SOTA) or comparable performance against existing feed-forward 3D reconstruction methods. To validate the effectiveness of our design, We also conduct several analyses: (1) a robustness evaluation against input image sequence permutations (Sec. 4.5), (2) an ablation study on scale-invariant point maps and affine-invariant camera poses (Sec. 4.6).
@@ 4.1 Camera Pose Estimation
Fast3R [42] sparse 0.095 0.065 0.144 0.089 0.673 0.759 0.135 0.091 0.163 0.104 0.759 0.877
CUT3R [36] 0.093 0.049 0.102 0.051 0.704 0.805 0.104 0.041 0.079 0.031 0.822 0.968
FLARE [47] 0.085 0.057 0.145 0.107 0.696 0.780 0.053 0.024 0.051 0.025 0.877 0.988
VGGT [34] 0.044 0.025 0.056 0.033 0.733 0.845 0.051 0.029 0.066 0.038 0.890 0.981
$\pi^{3}$ (Ours) 0.047 0.029 0.075 0.049 0.742 0.841 0.026 0.015 0.028 0.014 0.916 0.992
Fast3R [42] dense 0.040 0.017 0.056 0.018 0.644 0.725 0.072 0.030 0.050 0.016 0.790 0.934
CUT3R [36] 0.023 0.010 0.027 0.008 0.669 0.764 0.086 0.037 0.048 0.017 0.800 0.953
FLARE [47] 0.019 0.007 0.026 0.013 0.684 0.785 0.023 0.011 0.018 0.008 0.882 0.986
VGGT [34] 0.022 0.008 0.026 0.012 0.666 0.760 0.017 0.010 0.015 0.005 0.893 0.988
$\pi^{3}$ (Ours) 0.016 0.007 0.022 0.011 0.689 0.792 0.015 0.008 0.013 0.005 0.898 0.987
@@ 4.3 Video Depth Estimation
Abs Rel $\downarrow$ $\delta<1.25$ $\uparrow$ Abs Rel $\downarrow$ $\delta<1.25$ $\uparrow$ Abs Rel $\downarrow$ $\delta<1.25$ $\uparrow$
@@ 4.6 Ablation Study
Table 8: Ablation study on the key components of our model. We show how the performance metric improves as each component is added to the baseline.
Model ETH3D 7-Scenes NRGBD
Acc. $\downarrow$ Comp. $\downarrow$ N.C $\uparrow$ Acc. $\downarrow$ Comp. $\downarrow$ N.C $\uparrow$ Acc. $\downarrow$ Comp. $\downarrow$ N.C $\uparrow$
Mean Med. Mean Med. Mean Med. Mean Med. Mean Med. Mean Med. Mean Med. Mean Med. Mean Med.
Model 1 0.229 0.150 0.166 0.103 0.802 0.930 0.020 0.010 0.019 0.009 0.715 0.834 0.034 0.018 0.025 0.011 0.859 0.977
Model 2 0.197 0.118 0.118 0.065 0.820 0.943 0.020 0.009 0.020 0.008 0.716 0.837 0.031 0.018 0.023 0.010 0.861 0.978
Full Model 0.131 0.076 0.079 0.043 0.841 0.957 0.019 0.009 0.020 0.009 0.723 0.843 0.028 0.015 0.022 0.010 0.875 0.981
To validate the effectiveness of our proposed components, we conducted an ablation study by systematically removing features from our complete model. First, we created Model 2 by removing the affine-invariant camera pose modeling from our full model. Subsequently, we derived Model 1 by also removing the scale-invariant pointmap modeling from Model 2.
The primary difference between our full model and the ablated models (Model 1 and Model 2) is that the latter two incorporate a camera token. This token is essential for distinguishing the reference view, as the model is no longer permutation-equivariant after the removal of the affine-invariant camera pose modeling. At each iteration, the camera token is concatenated with a randomly selected reference view before the alternating-attention module similar to [34]. We compute an angle loss for rotation and a Huber loss for translation between the predicted and ground-truth poses in the reference view’s coordinate system for Model 1 and Model 2. While Model 1 and Model 2 share an identical architecture and parameter count, their key distinctions lie in the loss calculation and normalization processes. For Model 1, we neither perform alignment during the loss computation for the predicted pointmap nor do we normalize the pointmap itself. We found that applying normalization in this specific case led to anomalous and significantly degraded performance, a phenomenon also observed in prior work [34]. In contrast, the predicted local pointmaps are normalized for both Model 2 and the full model.
For a fair comparison, all models were trained for 80 epochs, with 800 iterations per epoch, on images with a resolution of $224\times 224$. They shared the same initialization procedure as our final model: we loaded pre-trained weights for the VGGT encoder and alternating-attention layers, and kept the encoder frozen throughout training. The comparative results for pointmap estimation across three datasets are presented in Table 8. For the 7-Scenes and NRGBD datasets, we use the same dense view setting as in the previous section.
We found that scale-invariant pointmap modeling does not yield significant performance gains on indoor datasets like 7-Scenes and NRGBD. For outdoor data, however, the performance improvement is substantially more pronounced. This observation is consistent with previous studies on scale-invariant depth, which have shown that outdoor scenes are more significantly affected by scale ambiguity. Furthermore, we observed that affine-invariant camera pose modeling consistently enhances the final performance. More importantly, unlike Model 1 and Model 2, its inclusion renders the model permutation-equivariant. Consequently, the model becomes robust to both the order of input frames and the selection of the reference view.
@@ 5 Conclusion
In this work, we introduced $\pi^{3}$, a feed-forward neural network that presents a new paradigm for visual geometry reconstruction by eliminating the reliance on a fixed reference view. By leveraging a fully permutation-equivariant architecture, our model is inherently robust to input ordering and leads to higher accuracy. This design choice removes a critical inductive bias found in previous methods, allowing our simple yet powerful approach to achieve state-of-the-art performance on a wide array of tasks, including camera pose estimation, depth estimation, and dense reconstruction. $\pi^{3}$ demonstrates that reference-free systems are not only viable but can lead to more stable and versatile 3D vision models.
@@ A.2 Training Details
We train $\pi^{3}$ in two stages, a process similar to Dust3R [39]. First, the model is trained on a low resolution of 224 $\times$ 224 pixels. Then, it is fine-tuned on images of random resolutions where the total pixel count is between 100,000 and 255,000 and the aspect ratio is sampled from the range [0.5, 2.0], a strategy similar to MoGe [37]. We use a dynamic batch sizing strategy similar to VGGT. In the first stage, we sample 64 images per GPU, and in the second stage, we sample 48 images per GPU. Each batch is composed of 2 to 24 images. Each training stage runs for 80 epochs, with each epoch comprising 800 iterations. Our final model is not trained from scratch. Instead, we initialize the weights for the encoder and the alternating attention module from the pre-trained VGGT model, and we keep the encoder frozen during training. We train the first stage on 16 A100 GPUs and the second stage on 64 A100 GPUs. For our loss function, we set the weights for each component as follows: $\lambda_{\text{normal}}=1.0$, $\lambda_{\text{conf}}=0.05$, $\lambda_{\text{cam}}=0.1$, and $\lambda_{\text{trans}}=100.0$. The implementation of our normal loss follows that of MoGe, and the resolution for aligning the local point map loss is set to 4096. Regarding optimization, we assign different initial learning rates to model components: $5\times 10^{-6}$ for the encoder and $5\times 10^{-5}$ for all other modules. We employ a OneCycleLR scheduler, where the learning rate anneals from its maximum value down to a minimal value over the entire training duration following a cosine curve. We use the same learning rate and scheduler settings for both stages. The confidence head is not trained jointly with the other modules. Instead, after completing the two main training stages, we freeze the rest of the network and train the confidence head in isolation. This final stage converges rapidly, typically within a few epochs, without impacting the model’s overall performance. We use gradient clipping with a norm of 1.0.